If two planes leave the same airport at 1:00 PM, how many miles apart will they be at 3:00 PM if one travels directly north at 150 mph and the other travels directly west at 200 mph?

Answer by checkley77(12844)

(2*150)^2+(2*200)^2=d^2
300^2+400^2=d^2
90,000+160,000=d^2
d^2=250,000
d=sqrt250,000
d=500 miles apart are the 2 planes after 2 hours.